OpenStudy (anonymous):

(5x + 10)^2 - 23 = 0

OpenStudy (anonymous):

I know I have to solve for x

OpenStudy (anonymous):

but I'm having trouble once I add 23

OpenStudy (jackellyn):

(5x+10)(5x+10)-23=0 FOIL and then find roots by either factoring or quadratic formula

OpenStudy (anonymous):

so far I have 25x + 50 + 50x + 100..is this correct so far?

OpenStudy (anonymous):

add like terms next right? so then I have 75x + 150..idk where to go from here

OpenStudy (jackellyn):

where is your x^2?

OpenStudy (jackellyn):

You should have \[25x ^{2}+100x+77=0\]

OpenStudy (anonymous):

so do I subtract 77 first?

OpenStudy (jackellyn):

From this point, you can tell that it is prime so you have to use the quadratic formula to solve for x. Do you know how to do that?

hero (hero):

I don't mean to interrupt, but @jackellyn, you're doing it the hard way.

OpenStudy (jackellyn):

Just doing it how I was taught @Hero but if you know a better way, go ahead

hero (hero):

\[(5x + 10)^2 - 23 = 0 \\(5x + 10)^2 = 23 \\5x+10 = \pm\sqrt{23} \\5x = \pm\sqrt{23} - 10 \\x = \frac{\pm\sqrt{23}-10}{5} \]

OpenStudy (jackellyn):

Did you understand what he did @mscrosscountry ?

OpenStudy (anonymous):

yes, this is how I was taught in class but I was just having trouble with what the next step would be once you square rooted. I understand now.